Aldridge, Inc. produces and sells a single product. Unit costs for direct materials and direct labor are

$24 and $36, respectively. In March, 20,000 units were produced, and total overhead was

$214,000. In April, 26,000 units were produced, and total overhead was $259,000. Selling price of

the unit is $100.

a. Compute the total fixed overhead of the company.

b. Compute the contribution margin ratio.

c. Compute the company's break-even point in sales dollars.

d. Compute the number of units which must be sold to have net income of $122,000 (ignore

income taxes).
Answer #1

a.

Variable cost per unit = Change in overhead /Change in output

= ( 259,000-214,000)/(26,000-20,000)

= 45,000/6,000

= $7.5

Total variable overheads at 20,000 units = 20,000 x 7.5

= $150,000

Fixed overheads at 20,000 units = Total overheads - Variable overheads

= 214,000-150,000

= $64,000

b.

Variable cost per unit = Direct materials + Direct labor + Variable overheads

= 24+36+7.5

= $67.5

Contribution margin per unit = Selling price per unit - Variable cost per unit

= 100-67.5

= $32.5

Contribution margin ratio = Contribution margin per unit/Selling price per unit

= 32.5/100

= 32.5%

c.

Break even sales in dollars = Fixed cost/Contribution margin ratio

= 64,000/32.5%

= $196,923

d.

Units to be sold to earn target profit = ( Fixed cost + Target profit)/Contribution margin per unit

= ( 64,000+122,000)/32.5

= 186,000/32.5

= 5,724 ( rounded to whole number)

Note: Exact answer may slightly differ due to rounding off

Kindly comment if you need further assistance. Thanks‼!